HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

COMPOSITE ABSORPTION CROSS SECTIONS AND INTEGRATED BAND INTENSITIES OF CHF_{3} IN SYNTHETIC AIR

\underline{K. BEREZKIN} , A. DOMANSKAYA, H. BOHLIUS, M.E. KIM, G. LI, and V. EBERT, Physikalisch-Technische Bundesanstalt (PTB), Bundesallee 100, 38116

Braunschweig, Germany

CHF_{3} (fluoroform, trifluoromethane, HFC-23, CAS no: 75-46-7) is a strong potent greenhouse gas (GWP-100 = 14600), with its atmospheric amount fraction increasing by 0.75 ± 0.02 pmol/mol/year^{1}. It is not only a byproduct of the synthesis of CHClF_{2} but is also used in the semiconductor industry and as a low-temperature refrigerant.

CHF_{3} is an oblate symmetric top molecule (C_{3v} point group) with three totally symmetric vibrations (ν_{1}, ν_{2}, ν_{3}) and three doubly degenerate vibrations (ν_{4}, ν_{5}, ν_{6}). Its congested absorption spectra contain not only transitions from ground state and hot bands for main isotopologue (^{12}CHF_{3}), but also information on ^{13}CHF_{3}, which bands appear nearby. Absorption in the region of ν_{2}/ν_{5} dyad occurs in the atmospheric transparency window and is used for remote sensing.

In the current work, we recorded several spectra with different compositions of CHF_{3} mixtures with synthetic air to reconstruct the cumulative spectrum at 296 K in the region of 550–2000 cm^{−1}. A similar algorithm was used in our work with CF_{4} and SF_{6} and in the PNNL database^{2}. CHF_{3} turned out to be a more challenging molecule since the contribution of the self-broadening becomes visible already at absorber dilution ratios exceeding 1%. This can lead to distortion (broadening) of the cumulative absorption cross sections compared to those corresponding to strongly diluted samples. Nevertheless, adding concentrated samples to the average does not affect the integrated intensities.

We report on the integrated intensities in the region of the ν_{3}, ν_{2}/ν_{5} and ν_{4} fundamental bands together with the overtone band of ν_{6} and two combinations that involve ν_{2} band, namely ν_{2} + ν_{6} and ν_{2} + ν_{3}. The expanded uncertainties (k = 2) are below 1.8% for the reported fundamental bands.

The measurements and analysis were performed within the frame of the EM-PIR project ”Metrology for Climate Relevant Volatile Organic Compounds” (Met-ClimVOC)^{3}. The spectra were recorded using a Bruker-125 HR spectrometer at the PTB EUMETRICSPEC infrastructure^{4}.

110.1016/j.jqsrt.2019.06.019, A. M. Fernando, P. F. Bernath, C. D. Boone, JQSRT, 238, 106540 (2019).

210.1366/0003702042641281, S.W. Sharpe et al., App. Spectrosc., 58(12), 1452–1461 (2004).

3EMPIR project ”Metrology for climate relevant volatile organic compounds”, (MetClimVOC).

4Spectral reference data for atmospheric monitoring, (EUMETRISPEC).

p-number: p243  Submitted on Tue, 27 Jun 2023 10:52:20 +0200